y'' = x + y2 + y'  y(0) = 1, y(1) = 0 calculate the differential equation using h = 1/3 with difference equations, it is enough to calculate y(1/3) and y(2/3).
Answer #1

# Griven y=re+y+y ycoged, 964) = 0, Using difference methood, with h= Y3 from , yco) = 0+ (13 °+ y60) = 5+ yco) agorin5 yil(l) = 2+ yCl). C Y!!)=0) How, y = 22 + y + - y(d) = 12 + (ych tych) = £t og to = 2 y!) = ] { y= 1+ y(t)=1t5 > 19(1)=